OpenStudy (anonymous):

I'm having a brain fart here. How many ml of 0.0825 M H2SO4 would you need to exactly neutralize 30.0 ml of 0.195 M NaOH? They want a factor label solution which i will put the one I came up with into the equation maker. But I had a prof tell me I was supose to do something else.

OpenStudy (anonymous):

\[30.0 ml NaOH(\frac{.195 mole NaOH}{1000 ml NaOH})(\frac{1 mole H_2SO_4}{1 mole NaOH})(\frac{1000 ml H_2SO_4}{.0825 mole H_2SO_4})=70.9 ml H_2SO_4\]

OpenStudy (anonymous):

This is the balanced equation that I used. \[NaOH + H_2SO_4 \rightarrow NaHSO_4 + H_2O\]

OpenStudy (jfraser):

the rxn should produce sodium sulfate, Na_2SO_4. Your mole factor should be\[(\frac{1mol H{_2}SO{_4}}{2mol NaOH})\] other than that you're alright

OpenStudy (xishem):

I have a question regarding neutralization. When you say that you are "neutralizing" a solution, are you essentially just wanting to react exactly all of the acid with a base (or vice-versa) to make non-acidic and non-basic products?

OpenStudy (jfraser):

yes.

OpenStudy (anonymous):

my numerical result is: 3.55mL H_2SO_4

OpenStudy (anonymous):

Oh well i guess i got that one wrong on the homework :(